Solve this application using logarithms.
At what interest rate compounded annually will money in savings double in five years?
I do not understand how to do the or what the logarithms mean in this problem. Thanks for your help

Answer by RAY100(1637)

Start with the compound interest formula
.
A = P (1+r) ^t,,,where A is Total, P is starting Principal,,,r is rate, t is time (yrs)
.
But to double means A =2P
.
2P = P(1+r)^t
.
divide by P thruout
.
2 = (1+r)^t
.
Take log both sides
.
log2 = log (1+r)^t = t log(1+r),,,, from log x^2=2logx
.
But,,,t=5,,,given
.
log2 =5log(1+r)
.
log2 /5 = .0602 = log(1+r)
.
Raise both sides to power of 10,,that is 10^.0602 =10^ (log (1+r))
.
1.148 = (1+r),,,remember 10^(log x) =x
.
(.148) =r,,,,,answer,,,or 14.8 %
.
checking
A/P = (1 +r)^5 = (1.148)^5 = 2,,,,ok